# Understanding Flatness in Generative Models: Its Role and Benefits

### v1

@@ Abstract
Flat minima, known to enhance generalization and robustness in supervised learning, remain largely unexplored in generative models. In this work, we systematically investigate the role of loss surface flatness in generative models, both theoretically and empirically, with a particular focus on diffusion models. We establish a theoretical claim that flatter minima improve robustness against perturbations in target prior distributions, leading to benefits such as reduced exposure bias—where errors in noise estimation accumulate over iterations—and significantly improved resilience to model quantization, preserving generative performance even under strong quantization constraints. We further observe that Sharpness-Aware Minimization (SAM), which explicitly controls the degree of flatness, effectively enhances flatness in diffusion models, whereas other well-known methods such as Stochastic Weight Averaging (SWA) and Exponential Moving Average (EMA), which promote flatness indirectly via ensembling, are less effective. Through extensive experiments on CIFAR-10, LSUN Tower, and FFHQ, we demonstrate that flat minima in diffusion models indeed improves not only generative performance but also robustness.
@@ 1 Introduction
What does it mean for a generative model to have a flat loss landscape? While flat minima have been extensively studied in supervised learning, where they are known to enhance generalization and robustness to distribution shifts [13, 5, 8, 17, 2, 28, 25], their role in generative models remains largely unexplored. At first glance, one might expect that flatness would be beneficial in generative models as well. However, upon deeper reflection, the connection is far less intuitive than in discriminative tasks. In supervised learning, flatter minima reduce sensitivity to small perturbations in the input space, thereby improving generalization. Yet, generative models—particularly those based on diffusion processes—operate on fundamentally different principles, as they take noise as input and iteratively refine it into structured outputs. This raises a fundamental question: “How does flatness affect generative modeling, and if beneficial, how can we induce it?”
@@
[…]
Table 1: Comparison of the 1) Frechet Inception Distance (FID) scores under 32- and 8-bit precision with (+/-) changes, 2) the gap of $\|\epsilon\|_{2}$ with training, 3) flatness metric called LPF for ADM and +SAM in the CIFAR-10 experiments. $\downarrow$: a lower value is preferred.
@@
As a preview of our analysis, Figures 1, 2, and Table 1 hint at the role and benefits of the generative models with flat minima. We find that SAM [8], an optimizer designed to promote flatter loss surfaces, significantly enhances flatness when applied to a baseline, i.e., Ablated Diffusion Model (ADM) [7], as shown in Figure 1 and Table 1 (measured using Low-Pass-Filter (LPF) flatness metric [2]). Regarding the benefits from flatness, we observe improved FID, reduced exposure biases (as indicated by the $\|\epsilon\|^{2}$ gap [31]), and lower degradation from model quantization (as reflected in 8-bit precision FID and qualitative results in Figure 2).
Returning to the key question, we set out to investigate the implication of flat loss landscapes in generative models, particularly diffusion models [12]. To make progress, we first seek to refine our core research question: “Why might flatter minima be desirable in generative models, and what theoretical guarantees can we establish?” To address these questions, we conduct a theoretical and empirical analysis of flat minima in diffusion models. First, we establish a theoretical link between the model’s parameter space and data probability density (Theorem 1), showing that flatter minima improve robustness to perturbations in the target prior distribution. As a consequence, we prove that the divergence between the true distribution and the learned distribution is upper-bounded, indicating that flat minima enable diffusion models to generalize not only to the training distribution but also to variations in perturbed data densities (Theorem 2). This insight suggests that flat diffusion models exhibit superior robustness when out-of-distribution is given, demonstrating their ability to cope with error accumulation problem during iterative sampling process, known as exposure bias problem [32, 31, 22]. Moreover, the loss flatness in the parameter space directly implies the robustness against model parameter changes, which indicates that the flatness offers the robustness of generative performance when the model has been quantized.
@@
To validate these findings, we conduct extensive experiments on CIFAR-10, LSUN Tower, and FFHQ. Our results show that flat minima in diffusion models consistently improve generative performance with enhanced generalization. This improved generalization offers several practical advantages. For instance, we observe that flat diffusion models better mitigate exposure bias, where errors in noise estimation accumulate over iterations, leading to more stable sampling dynamics. We also show that flat diffusion models results a low quantization error where the model precision is reduced from 32-bit to 8-bit.
@@
  • We empirically analyze SWA [13] and EMA, showing that while they improve FID, they do not significantly enhance flatness. In contrast, SAM reshapes the loss surface and effectively promotes flatness in diffusion models.
@@ 2.1 Optimization for flatness
8:        $w\leftarrow w-\alpha\nabla\mathcal{L}_{i}(\hat{w})$Descent step}
@@ 2.2 Score-based Generative Models
where $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ is the time-dependent score function, approximated by $s_{\theta}(\mathbf{x},t)$, and $d\bar{\mathbf{w}}$ is a standard Wiener process running in reverse time. To successfully estimate the score function, score matching objective is defined as:
@@
2.3 Diffusion models
SGMs are closely related to diffusion probabilistic models (DPMs) [37, 12], which DPMs can be viewed as a discrete-time approximation of SGM with following forward diffusion process:
@@ 2.3 Diffusion models
where $\beta_{t}$ is a predefined parameter by variance scheduling, $\mathcal{N}(m;\sigma^{2})$ is a Guassian distribution with mean $m$ and variance $\sigma^{2}$, and $\mathbf{I}$ is an identity matrix. Given an initial sample $\mathbf{x}_{0}$, the noisy sample at timestep $t$ is generated as follows:
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}(0,\mathbf{I})$ (6)
where $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$.
Then, we can analytically derive the posterior distribution of injected gaussian noise as follows:
@@
To estimate the posterior distribution, diffusion models parameterize $\tilde{\mu}(\mathbf{x}_{t},\mathbf{x}_{0})$ or instead $\epsilon_{t}$ after applying reparameterization trick. Once trainig is over, diffusion model produce the clean images by iteratively removing the gaussian noise as follows:
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t))+\beta_{t}\mathbf{z},.$ (10)
where $\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$.
While these iterative denoising frameworks successfully learn the underlying distribution and synthesize images with exceptional fidelity, previous works [32, 31, 22] have raised an input discrepancy problem between the training and sampling phases, known as exposure bias problem. Specifically, during training, the diffusion model is provided with ground truth noisy images (Equation (6)), whereas during sampling, it receives a denoised image (Equation (10)) came from the previous timestep. The presence of errors in these inputs makes it challenging for the diffusion model to accurately predict noise. As these errors accumulate over iterative timesteps, they cause the generated samples to significantly deviate from the desired trajectory.
@@ 3.1 Notations
Our analysis is built upon the settings of SGM, which is a fundamental form of diffusion models. Beyond the brief preliminaries in Section 2.2, we introduce additional notations and formulations to be used in our analysis. By following [23], we formulate the score model $s_{\boldsymbol{\theta}}(\cdot,\cdot)$ as a random feature model:
$s_{\boldsymbol{\theta}}(\mathbf{x},t):=\frac{1}{m}\mathbf{\boldsymbol{\theta}}\sigma\big{(}\mathbf{W}^{T}\mathbf{x}+\mathbf{U}^{T}\mathbf{e}_{t}\big{)},$ (11)
where $\mathbf{x}\in\mathbb{R}^{d\times 1}$ is the given input, $\mathbf{e}_{t}$ is the timestep embedding for $t$, a group of parameters includes $\boldsymbol{\theta}\in\mathbb{R}^{d\times m}$, $\mathbf{W}\in\mathbb{R}^{d\times m}$ and $\mathbf{U}\in\mathbb{R}^{d_{e}\times m}$, and a few positive integers include $d$, $m$, and $d_{e}$. By following the training procedures of diffusion models, $\boldsymbol{\theta}$ is the learnable parameter, while others, i.e., $\mathbf{W}$ and $\mathbf{U}$, which respectively embed $\mathbf{x}$ and $\mathbf{e}_{t}$, are set to be frozen. For a given timestep $t$, the score matching loss objective defined in Equation 4
@@
where $p_{t}(\mathbf{x})$ is the probability density function for the prior distribution of $\mathbf{x}$.
@@ 3.2 Mathematical Claims
For the main claims, we follow the formulation of Equation (12), while omitting timestep $t$ without loss of generality. Our mathematical claims are valid for all timesteps. We start from defining flat minima in SGMs (Definition 1), and the robustness against the gap of ground truth prior $p(\mathbf{x})$ and the estimated prior $\hat{p}(\mathbf{x})$ (Definition 2). For the flat minima definition, we rehearse the definition of flat minima for the supervised learning case [36], thus $\Delta$-flat minima of SGMs with the score matching loss is defined as follows:
@@
where $l^{*}:=\mathcal{L}(\mathbf{x};\boldsymbol{\theta}^{*},p)$ and $\Delta\in\mathbb{R}^{+}$.111$\forall$ means “for all,” $\exists$ means “there exists,” and $\mathbb{R}^{+}$ indicates the set of positive real numbers.
@@
$\mathcal{L}(\mathbf{x};\boldsymbol{\theta}+\delta,p)=\mathcal{L}(\mathbf{x};\boldsymbol{\theta},\hat{p}).$ (13)
This means how the parameter perturbation is translated into the shift of the prior, while keeping the loss unchanged. Let us provide a mathematical claim of it:
@@
[…]
This theorem shows that the parameter perturbation can be translated into the “scaling” of probability density function. However, the scaling depends on various parameters, thus it is nontrival to imagine the behavior of $\hat{p}$. Let us narrow our view to the diffusion model, where it predicts a noise signal sampled from Gaussian distribution, i.e., $\mathcal{N}(0,\mathbf{I})$. It changes our focus on handling the noise distribution, i.e., $\epsilon$, rather than the distribution of data side, i.e., $p$. The following corollary describes how the perturbed $\hat{\epsilon}$ looks like:
@@
Next, let us consider all possible perturbations within the ball of norm, i.e., $\|\Delta\|_{2}\leq\Delta$. The resulting set of perturbed distributions induced by the parameter perturbation is defined as follows:
@@
(A set of perturbed distribution) For a given distribution of $p(\mathbf{x})$, a set of distributions $\hat{\mathcal{P}}(\mathbf{x};p,\Delta)$ is defined as the set of perturbed distributions $\hat{p}(\mathbf{x})$:
@@
(A link from $\Delta$-flatness to $\hat{\mathcal{P}}$) A $\Delta$-flat minimum $\boldsymbol{\theta}^{*}$ achieves the flat loss values for all distributions sampled from the set of perturbed distribution:
@@
We finally link flatness to the distributional gap, i.e., connecting Definitions 1 and 2. Let us anticipate the distribution in $\hat{\mathcal{P}}$ with the maximal divergence from $p$. The distribution gap $\mathcal{E}$ is then upper bounded by the divergence of the most outreached distribution:
Figure 3: A conceptual illustration of theoretical analysis. Theorem 1 (Corollary 1 for diffusion model) translates the perturbation in the parameter space into the set of perturbed distributions. Theorem 2 (Corollary 2 for diffusion model) shows that flat minima leads to the robustness against the distribution gap.
@@
+SAM 9.01 3.83 16.02 4.79 31.89 12.43
+SAM+EMA 7.00 3.18 6.66 2.30 21.48 6.44
+SAM+SWA 7.27 2.96 6.50 2.27 21.11 5.32
Table 2: FID Scores for ADM baselines with different algorithms on CIRAR-10, LSUN Tower, and FFHQ dataset. We use DDPM sampling with shorter respacing timesteps, $T^{\prime}=20,100$.
@@
$\mathcal{E}\leq\operatorname*{arg\,max}_{\hat{p}\sim\hat{\mathcal{P}}(\mathbf{x};p,\Delta)}D(p||\hat{p}).$ (19)
Let us further manipulate the upper bound for the diffusion model case. From Corollary 1, $\hat{p}$ is a form of Gaussian distribution; thus it is possible to achieve the closed form of the upperbound as follows:
@@
$\mathcal{E}$ $\leq\operatorname*{arg\,max}_{\|\delta\|_{2}\leq\Delta}\frac{1}{2}\Bigg{[}\log|\Sigma_{\delta}|-d+tr(\Sigma_{\delta}^{-1})+{\boldsymbol{\mu}_{\delta}}^{T}{\Sigma_{\delta}}^{-1}\boldsymbol{\mu}_{\delta}\Bigg{]}$
$\leq\frac{1}{2}\Bigg{[}\sum_{i=1}^{d}\log\frac{m}{m+\sigma_{i}}+\sum_{i=1}^{d}\frac{\sigma_{i}}{m}+\frac{\sigma_{d}}{m^{2}}\|\mathbf{U}^{T}\mathbf{e}\|_{2}^{2}\Delta^{2}\Bigg{]},$ (20)
@@
(Flatter minima enhance distribution gap robustness) It is noteworthy that flatter minima with a large $\Delta$ flatten the loss values for the far-pushed away perturbed distribution $\hat{p}$ with large divergence, leading to the larger bound of $\mathcal{E}$. It indicates that flattening the loss surface makes the generative model zero-force the loss values of diversified prior distribution, thus enhancing the robustness against the prior distribution gap. On the other hand, $m$ is the number of trainable parameters, which is quite large, so that the first and second terms are negligible.
Flatness reduces exposure bias (a gap of $\hat{\epsilon}$ and $\epsilon$): A clear advantage of flatness is robustness to exposure bias, where the errors in noise estimation accumulate over iterations, thus severely deteriorates the generative performance. In formula, $\hat{\epsilon}$ deviates from $\epsilon$. We argue that a generative model with a flat minimum suppresses the loss values of perturbed estimation, so that the error accumulation is sufficiently relieved. To be shown in experiments, we empirically confirm that flat minima tend to show a smaller exposure bias, leading to the robust generative performance.
Flatness becomes robust to model quantization (a compression from $\boldsymbol{\theta}$ to $\hat{\boldsymbol{\theta}}$): For another benefit, flat minima make the model robust to the performance degradation caused by the model quantization. Recently, quantization of generative models becomes crucial, when enabling real-time applications of generative tasks [44, 26, 27]. We claim that flat minima inherently bolster the robustness against the degradation due to the quantization, because the quantized model, $\hat{\boldsymbol{\theta}}$, can be viewed as a perturbed version of the full-precision parameter, i.e., $\hat{\boldsymbol{\theta}}=\boldsymbol{\theta}+\Delta$; thus the loss remains flat for the perturbation [27]. In experiments, we clearly demonstrate that the diffusion model with flatness is remarkably resilient to the degradation from quantization than other methods.
A conceptual overview of our theoretical analysis is illustrated in Figure 3. The proof of Theorem 1, and Corollary 1, 2 are given in Appendix 1 in Supplementary.
Figure 4: Generation results of (a) full precision (32-bit) and (b) 8-bit quantized models. We use respaced timesteps $T^{\prime}=20$ for sampling. For the 8-bit case, we selectively compare +EMA, +SAM and their combinations, where the relatively flatter minima have been found.

### v2

@@ Abstract
Flat minima, known to enhance generalization and robustness in supervised learning, remain largely unexplored in generative models. In this work, we systematically investigate the role of loss surface flatness in generative models, both theoretically and empirically, with a particular focus on diffusion models. We establish a theoretical claim that flatter minima improve robustness against perturbations in target prior distributions, leading to benefits such as reduced exposure bias—where errors in noise estimation accumulate over iterations—and significantly improved resilience to model quantization, preserving generative performance even under strong quantization constraints. We further observe that Sharpness-Aware Minimization (SAM), which explicitly controls the degree of flatness, effectively enhances flatness in diffusion models even surpassing the indirectly promoting flatness methods—Input Perturbation (IP) which enforces the Lipschitz condition, ensembling-based approach like Stochastic Weight Averaging (SWA) and Exponential Moving Average (EMA)—are less effective. Through extensive experiments on CIFAR-10, LSUN Tower, and FFHQ, we demonstrate that flat minima in diffusion models indeed improve not only generative performance but also robustness.
@@ 1 Introduction
What does it mean for a generative model to have a flat loss landscape? While flat minima have been extensively studied in supervised learning, where they are known to enhance generalization and robustness to distribution shifts, e.g., promoting stable label predictions under input shifts such as domain changes [14, 5, 9, 18, 2, 27, 24], their role in generative models remains largely unexplored.
@@
ADM 34.47 $\overset{\text{{\tiny{\color[rgb]{1,0,0}+13.65}}}}{\longrightarrow}$ 48.02 +11.39 0.097
+SAM 9.01 -$\overset{\text{{\tiny{\color[rgb]{0,0.6,0}-0.07}}}}{\longrightarrow}$ 8.94 +3.32 0.063
Table 1: Comparison of the 1) Fréchet Inception Distance (FID) scores under 32- and 8-bit precision with (+/-) changes, 2) the gap of $\|\epsilon\|_{2}$ with training, 3) flatness metric called LPF for ADM and +SAM in the CIFAR-10 experiments. $\downarrow$: a lower value is preferred.
@@
At first glance, one might expect that flatness would be beneficial in generative models as well. However, generative models—particularly those based on diffusion processes—operate under fundamentally different principles: rather than mapping structured inputs to labels, they take noise as input and iteratively refine it into coherent outputs through a learned denoising process. This key difference motivates a critical question: Do flat minima compel the generative model to produce similar contents regardless of input variations? If so, such behavior would contradict the goal of generative diversity and potentially impair generalization. If not, it remains unclear how flatness affects generative modeling, and if it is indeed beneficial, how can we induce it?
As a preview of our analysis, Figure 1, 2 and Table 1 hint at the role and benefits of the generative models with flat minima. We find that SAM [9], an optimizer designed to promote flatter loss surfaces, significantly enhances flatness when applied to a baseline, i.e., Ablated Diffusion Model (ADM) [7], as shown in Figs. 1 and 1 (measured using Low-Pass-Filter (LPF) flatness metric [2]). Regarding the benefits from flatness, we observe improved FID, reduced exposure biases (as indicated by the $\|\epsilon\|^{2}$ gap [31]), and lower degradation from model quantization (as reflected in 8-bit precision FID and qualitative results in Fig. 2).
Returning to the key question, we set out to investigate the implication of flat loss landscapes in generative models, particularly diffusion models [13]. To make progress, we first seek to refine our core research question: “Why might flatter minima be desirable in generative models, and what theoretical guarantees can we establish?”
To address these questions, we conduct a theoretical and empirical analysis of flat minima in diffusion models. First, we establish a theoretical link between the model’s parameter space and data probability density (Theorem 1), showing that flatter minima improve robustness to perturbations in the target prior distribution. As a consequence, we prove that the divergence between the true distribution and the learned distribution is upper-bounded, indicating that flat minima enable diffusion models to generalize not only to the training distribution but also to variations in perturbed data densities (Theorem 2). This insight suggests that flat diffusion models exhibit superior robustness when out-of-distribution is given, demonstrating their ability to cope with the error accumulation problem during the iterative sampling process, known as exposure bias problem [32, 31, 21]. Moreover, the flat loss landscape in the parameter space directly implies the robustness against model parameter changes, which indicates that the flatness offers the robustness of generative performance when the model has been quantized.
@@
To validate these findings, we conduct extensive experiments on CIFAR-10, LSUN Tower, and FFHQ. Our results show that flat minima in diffusion models consistently improve generative performance with enhanced generalization. This improved generalization offers several practical advantages. For instance, we observe that flat diffusion models better mitigate exposure bias, where errors in noise estimation accumulate over iterations, leading to more stable sampling dynamics. We also show that flat diffusion models result in a low quantization error where the model precision is reduced from 32-bit to 8-bit.
@@
  • We empirically analyze IP [32], SWA [14], and EMA, showing that while they improve FID, they do not significantly enhance flatness. In contrast, SAM reshapes the loss surface and effectively promotes flatness in diffusion models.
@@ 2.1 Optimization for flatness
[…]
@@
IP. IP [32] was introduced to address the exposure bias problem, which inherently arises in the auto-regressive denoising process of diffusion models. They define exposure bias as a generalization failure to unseen inputs, caused by the accumulation of prediction errors over timesteps. To mitigate this issue, authors propose a smoothed diffusion model that satisfies a Lipschitz condition by training on randomly perturbed input data. This strategy aligns with the flat minima perspective, implicitly promoting generalization to nearby unseen inputs.
@@ 2.2 Score-based Generative Models
where $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ is the time-dependent score function, approximated by $s_{\theta}(\mathbf{x},t)$, and $d\bar{\mathbf{w}}$ is a standard Wiener process running in reverse time. To successfully estimate the score function, the score matching objective is defined as:
@@
Diffusion models. Diffusion models (DPMs) [38, 13] can be viewed as a discrete-time approximation of SGMs, with the following forward and backward diffusion process:
@@
where $\beta_{t}$ is a predefined parameter by variance scheduling, $\alpha_{t}:=1-\beta_{t}$, and $\bar{\alpha}_{t}:=\Pi_{s=1}^{t}\alpha_{s}$.
Exposure bias. Previous works [32, 31, 21] have raised an input discrepancy problem between the training and sampling phases, known as exposure bias problem. Specifically, during training, the DPM is provided with ground truth noisy images, whereas during sampling, it receives a denoised image that came from the previous timestep. The presence of errors in these inputs, amplified through the iterative denoising process, makes it challenging for the diffusion model to accurately predict noise. As these errors accumulate over iterative timesteps, they cause the generated samples to significantly deviate from the desired trajectory.
@@ 3.1 Notations
Our analysis is built upon the settings of SGM, which is a fundamental form of diffusion models. Beyond the brief preliminaries in Section 2.2, we introduce additional notations and formulations to be used in our analysis.
Following prior theoretical analysis [22], we formulate the simplified score model $s_{\theta}(\cdot,\cdot)$ as a random feature model consisting of an encoder and a decoder.
$s_{\boldsymbol{\theta}}(\mathbf{x},t):=\frac{1}{m}\mathbf{\boldsymbol{\theta}}\sigma\big{(}\mathbf{W}^{\top}\mathbf{x}+\mathbf{U}^{\top}\mathbf{e}_{t}\big{)},$ (9)
where $\mathbf{x}\in\mathbb{R}^{d\times 1}$ is the given input, $\mathbf{e}_{t}$ is the timestep embedding for $t$, a group of parameters includes $\boldsymbol{\theta}\in\mathbb{R}^{d\times m}$, $\mathbf{W}\in\mathbb{R}^{d\times m}$ and $\mathbf{U}\in\mathbb{R}^{d_{e}\times m}$, and a few positive integers include $d$, $m$, and $d_{e}$. By following the training procedures of diffusion models, $\boldsymbol{\theta}$ is the learnable parameter, while others, i.e., $\mathbf{W}$ and $\mathbf{U}$, which respectively embed $\mathbf{x}$ and $\mathbf{e}_{t}$, are set to be frozen.
For a given timestep $t$, the score matching loss objective defined in Eq. 4
@@
where $p_{t}(\mathbf{x})$ is the probability density function for the target distribution at time $t$.
@@ 3.2 Mathematical Claims
For the main claims, we follow the formulation of Equation (10), while omitting timestep $t$ without loss of generality. Our mathematical claims are valid for all timesteps. We start from defining flat minima in SGMs (Definition 1), and the robustness against the gap of ground truth prior $p(\mathbf{x})$ and the estimated prior $\hat{p}(\mathbf{x})$ (Definition 2). For the flat minima definition, we rehearse the definition of flat minima for the supervised learning case [37] and extend it to the $\Delta$-flat minima of SGMs.
@@
[…]
@@
[…]
This means how the parameter perturbation is translated into the shift of the prior, while keeping the loss unchanged. Let us provide a mathematical claim for it:
@@
where $I(\mathbf{x},\delta):=\frac{1}{2}\mathbf{x}^{\top}(\delta\mathbf{W}^{\top})\mathbf{x}+\mathbf{x}^{\top}\delta(\mathbf{U}^{\top}\mathbf{e})+C$ with $\delta\mathbf{W}^{\top}$ being symmetric, and $C\in\mathbb{R}$ is set to satisfy $\int_{\mathbb{R}^{d}}e^{-I(\mathbf{x},\delta)}p(\mathbf{x})d\mathbf{x}=1$.
For analytical convenience, Theorem 1 derives a perturbed density under symmetric $\delta\mathbf{W}^{\top}$, showing that parameter perturbation translates into a “scaling” of probability density. Note that $C$ absorbs the normalizing constant. However, the scaling depends on various parameters. Thus, it is non-trivial to imagine the behavior of $\hat{p}$. Let us narrow our view to the diffusion model, where it predicts a noise signal sampled from a Gaussian distribution, i.e., $\mathcal{N}(0,\mathbf{I})$. It changes our focus on handling the noise distribution, i.e., $\epsilon$, rather than the distribution of data side, i.e., $p$. The following corollary describes how the perturbed $\hat{\epsilon}$ looks like:
@@
where $\Sigma_{\delta}:=\bigg{(}\mathbf{I}+\displaystyle\frac{\delta\mathbf{W}^{\top}}{m}\bigg{)}^{-1}$, $\boldsymbol{\mu}_{\delta}:=\displaystyle\frac{1}{m}\Sigma_{\delta}\delta\mathbf{U}^{\top}\mathbf{e}$.
@@
Next, let us consider all possible perturbations within the ball of norm, i.e., $\|\delta\|_{2}\leq\Delta$ and $\delta\mathbf{W}^{\top}=\mathbf{W}\delta^{\top}$. The resulting set of perturbed distributions induced by the parameter perturbation is defined as follows:
@@
(A set of perturbed distributions) For a given distribution of $p(\mathbf{x})$, a set of distributions $\hat{\mathcal{P}}(\mathbf{x};p,\Delta)$ is defined as the set of perturbed distributions $\hat{p}(\mathbf{x})$:
$\hat{\mathcal{P}}$ $(\mathbf{x};p,\Delta)$
$:=\left\{e^{-I(\mathbf{x},\delta)}p(\mathbf{x})\middle|\|\delta\|_{2}\leq\Delta,\delta\mathbf{W}^{\top}=\mathbf{W}\delta^{\top}\right\}.$ (14)
@@
(A link from $\Delta$-flatness to $\hat{\mathcal{P}}$) A $\Delta$-flat minimum $\boldsymbol{\theta}^{*}$ achieves the flat loss values for all distributions sampled from the set of perturbed distributions:
@@
We finally link flatness to the distributional gap, i.e., connecting Definition 1 and 2. Let us anticipate the distribution in $\hat{\mathcal{P}}$ with the maximal divergence from $p$. The distribution gap $\mathcal{E}$ is then upper bounded by the divergence of the most outreached distribution:
Figure 3: A conceptual illustration of theoretical analysis. Theorem 1 (Corollary 1 for diffusion model) translates the perturbation in the parameter space into the set of perturbed distributions. Theorem 2 (Corollary 2 for diffusion model) shows that flat minima lead to robustness against the distribution gap.
@@
+IP 20.11 7.23 25.77 7.00 15.03 13.55
+IP+EMA 9.10 3.46 7.66 2.43 11.72 4.00
+IP+SWA 9.04 3.07 8.55 2.34 12.99 3.54
+SAM 9.01 3.83 16.02 4.79 11.59 5.29
+SAM+EMA 7.00 3.18 6.66 2.30 11.41 5.04
+SAM+SWA 7.27 2.96 6.50 2.27 12.15 4.17
Table 2: FID Scores for ADM baselines with different algorithms on CIRAR-10, LSUN Tower, and FFHQ datasets. We use DDPM sampling with shorter respacing timesteps, $T^{\prime}=20,100$.
@@
[…]
Let us further manipulate the upper bound for the diffusion model case. From Corollary 1, $\hat{p}$ is a form of Gaussian distribution; thus, it is possible to achieve the closed form of the upper bound as follows:
@@
$\mathcal{E}$ $\leq\max_{\|\delta\|_{2}\leq\Delta}\frac{1}{2}\Bigg{[}\log|\Sigma_{\delta}|-d+tr(\Sigma_{\delta}^{-1})+{\boldsymbol{\mu}_{\delta}}^{\top}{\Sigma_{\delta}}^{-1}\boldsymbol{\mu}_{\delta}\Bigg{]}$
$\leq\frac{1}{2}\Bigg{[}\sum_{i=1}^{d}(\sigma_{i}-\log\sigma_{i})-d+\frac{\sigma_{d}}{m^{2}}\|\mathbf{U}^{\top}\mathbf{e}\|_{2}^{2}\Delta^{2}\Bigg{]},$ (18)
@@
(Flatter minima enhance distribution gap robustness) It is noteworthy that flatter minima with a large $\Delta$ flatten the loss values for the far-pushed away perturbed distribution $\hat{p}$ with large divergence, leading to the larger bound of $\mathcal{E}$. It indicates that flattening the loss surface makes the generative model zero-force the loss values of diversified prior distribution, thus enhancing the robustness against the prior distribution gap.
Flatness reduces exposure bias (a gap of $\hat{\epsilon}$ and $\epsilon$): A clear advantage of flatness is robustness to exposure bias, where the errors in noise estimation accumulate over iterations, thus severely deteriorating the generative performance. In formula, $\hat{\epsilon}$ deviates from $\epsilon$. We argue that a generative model with a flat minimum suppresses the loss values of perturbed estimation, so that the error accumulation is sufficiently relieved. To be shown in experiments, we empirically confirm that flat minima tend to show a smaller exposure bias, leading to robust generative performance.
Flatness becomes robust to model quantization (a compression from $\boldsymbol{\theta}$ to $\hat{\boldsymbol{\theta}}$): For another benefit, flat minima make the model robust to the performance degradation caused by the model quantization. Recently, the quantization of generative models has become crucial, when enabling real-time applications of generative tasks [45, 25, 26]. We claim that flat minima inherently bolster the robustness against the degradation due to the quantization, because the quantized model, $\hat{\boldsymbol{\theta}}$, can be viewed as a perturbed version of the full-precision parameter, i.e., $\hat{\boldsymbol{\theta}}=\boldsymbol{\theta}+\Delta$; thus the loss remains flat for the perturbation [26]. In experiments, we clearly demonstrate that the diffusion model with flatness is remarkably resilient to degradation from quantization compared to other methods.
A conceptual overview of our theoretical analysis is illustrated in Fig. 3. The proof of Theorem 1, and Corollary 1, 2 are given in Appendix A in Supplementary.